OpenStudy (anonymous):

Supplementary angles are 2 angles whose sum is 180°. ∠A and ∠B are supplementary angles. m∠A = 5x - 5 and m∠B = 4x + 20.

OpenStudy (anonymous):

what does angle a and b equal to?

OpenStudy (anonymous):

(5x - 5) + (4x + 20) = 180 5x - 5 + 4x + 20 = 180 9x + 15 = 180 9x + 15 - 15 = 180 - 15 9x = 165 9x/9 = 165/9 x = 18.33333333333333333 A = 5x - 5 = 5(18.333333333) - 5 = 86.666666666666666 B = 4x + 20 = 4(18.3333333333) + 20 = 93.33333333333 ~Approximate answers~